Natalie mixed 11 oz. of peanuts with 9 oz. of mixed nuts containing 60% peanuts. Peanuts are what percent of the new mixture?
What do x and y represent?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Natalie mixed 11 oz. of peanuts with 9 oz. of mixed nuts containing 60% peanuts. Peanuts are what percent of the new mixture?
-----
peanuts + peanuts = peanuts
11 + 0.60*9 = x(11+9)
11 + 5.4 = 20x
16.4 = 20x
x = 0.82 or 82% (percent of peanuts in the final mixture)
======================
Cheers,
Stan H.
=============

Answer by mananth(16946)

--------percent ---------------- quantity
peanuts 100 ---------------- 11 Oz
mixed nuts 60 ---------------- x Oz
Mixture 90 ---------------- 11 + x Oz

100*11+ 60x=90(11+x)
1100 +60x =990 +90x
60x+ -90x = 990- -1100
-30 x = -110
/ -30
x = 3.67
3.67 Oz of mixed nuts
Peanuts = (11-3.67)= 7.33 Oz
Percentage of peanuts = ((7.33)/(14.67)) * 100 = 49.97